Input


The first number, second number, third number, fourth number and fifth number contains:


Description

N/A


Constraints

This may be a decimal number.


Sample

N/A


Output


The first line will contain a message prompt to input the first number.

The second line will contain a message prompt to input the second number.

The third line will contain a message prompt to input the third number.

The fourth line will contain a message prompt to input the fourth number.

The fifth line will contain a message prompt to input the fifth number.

The last line contains the sum of all negative numbers rounded off and displayed with 2 decimal places.


Enter·the·first·number:·34

Enter·the·second·number:·43.5

Enter·the·third·number:·-53.24

Enter·the·fourth·number:·56.3

Enter·the·fifth·number:·-10.32

Sum·of·all·negative·integers·=·-63.56


1
Expert's answer
2021-09-11T07:19:38-0400
F1=float(input("Enter the first number: "))
F2=float(input("Enter the second number: "))
F3=float(input("Enter the third number: "))
F4=float(input("Enter the fourth number: "))
F5=float(input("Enter the fifth number: "))
if F1<0:
    F1=F1
else:
    F1=0
if F2<0:
    F2=F1+F2
else:
    F2=F1
if F3<0:
    F3=F2+F3
else:
    F3=F2
if F4<0:
    F4=F4+F3
else:
    F4=F3
if F5<0:
    F5=F5+F4
else:
    F5=F4
print("Sum of all negative integers=",F5)